# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v2

@@ 3.2.8. The moment maps $\boldsymbol{m}$ and $\boldsymbol{m}_{N}$
where $h_{W}$ denotes the Hamiltonian of a real vector field $W$ on $X$ corresponding to an element $W$ in $\mathfrak{k},$ defined by $h_{W}\coloneqq\mathcal{L}_{JW}\phi$ (defined as in formula 3.18). We will denote by $\boldsymbol{m}_{N}$ the moment map of the diagonal action of the group $\mathcal{K}$ on $X^{N},$ endowed with the symplectic form
@@
Moreover, 0 is a regular value of $\boldsymbol{m}_{N}$ iff all the stabilizers in $\mathcal{K}$ of the points in $\left\{\boldsymbol{m}_{N}=0\right\}$ are finite iff the semistable locus of $X^{N}$ coincides with the polystable locus [64, Cor 3.10].
@@ 3.3. The structure of the space $\mathcal{P}(X)_{0}$ of probability measures with vanishing moment
Consider the infimum of the pluricomplex energy $E$ on the orbit $\mathcal{G}\nu$ of any volume form in $\mathcal{P}(X)$ with divisorial singularities along $\Delta.$ It is attained at a unique $\mu$ in $(\mathcal{G}\nu)/\mathcal{K},$ which is the unique critical point of $E$ on $(\mathcal{G}\nu)/\mathcal{K}.$ Moreover, if the Futaki character of $(X,\Delta)$ vanishes, then $\mu$ is the unique element in $\left(\mathcal{G}\nu\cap\mathcal{P}(X)_{0}\right)/\mathcal{K}$ and, more generally, any $\mu\in\mathcal{P}(X)_{0}$ minimizes $E$ on $\mathcal{G}\mu.$
@@
First note that $E$ is $\mathcal{K}$–invariant (since $\omega_{0}$ is $\mathcal{K}$–invariant). We thus obtain a function
$\mathcal{G}/\mathcal{K}\rightarrow\mathbb{R},\,\,\,g\mapsto E(g^{*}\nu),$
The first part of the proposition to be proved says that this function has a unique critical point. To prove this first recall that $\mathcal{G}=\mathcal{K}\exp(J\mathfrak{k}),$ by polar decomposition of complex reductive Lie groups. Hence, by basic convex analysis it enough to show that $E(g^{*}\nu)$ is proper and strictly convex along curves $g_{t}$ of the form $g_{t}=\exp(tJW),$ where $W\in\mathfrak{k}-\{0\}.$ Expressing $\nu=(dd^{c}\phi_{\nu})^{n}/V,$ for some $\phi_{\nu}\in\text{PSH}(-K_{(X,\Delta)})_{b},$ gives $g_{t}^{*}\nu=dd^{c}(g_{t}^{*}\phi_{\nu}))^{n}/V.$ However, the real one-parameter curve $g_{t}^{*}\phi_{\nu}$ need not define a geodesic in the sense of [40], even when $\phi_{\nu}$ is smooth, unless $\phi_{\nu}$ is $\mathcal{K}$-invariant. To bypass this difficulty we will use a complexification argument, by setting $\tau=s+it$ for $s,t\in\mathbb{R}$ and $g_{\tau}\coloneqq\exp(sW+tJW).$ Since $E$ is $\mathcal{K}$-invariant we have $E(g_{{}_{\tau}}^{*}\nu)=E(g_{t}^{*}\nu).$ The corresponding properness on $\mathcal{G}/\mathcal{K}$ is well-known (see, for example, [21, Remark 9.2]). What remains is thus to show that
(3.23) $\partial_{\tau}\bar{\partial}_{\tau}E(g_{{}_{\tau}}^{*}\nu)>0.$
[…]
$\phi^{\tau}\coloneqq\exp(g_{\tau})^{*}\phi_{\nu},$
we can express $g_{\tau}^{*}\nu=(dd^{c}\phi^{\tau})^{n}/V.$ We claim that there exists a positive constant $c$ such that
(3.24) $\frac{\partial}{\partial\tau}\frac{\partial}{\partial\overline{\tau}}E\left(\frac{(dd^{c}\phi^{\tau})^{n}}{V}\right)\geq c\int_{X}i\partial\left(\frac{\partial\phi^{\tau}}{\partial\overline{\tau}}\right)\wedge\overline{\partial\left(\frac{\partial\phi^{\tau}}{\partial\overline{\tau}}\right)}\wedge dd^{c}\phi_{0}\wedge(dd^{c}\phi^{\tau})^{n-2}$
(when $n=2$ and when $n=1$ the last factor above should be omitted). Accepting this for the moment, it follows that $\partial_{\tau}\bar{\partial}_{\tau}E(g_{{}_{\tau}}^{*}\nu)\geq 0.$ Moreover, if $\partial_{\tau}\bar{\partial}_{\tau}E(g_{{}_{\tau}}^{*}\nu)=0,$ then $\partial(\frac{\partial\phi^{\tau}}{\partial\overline{\tau}})=0$ on $X-\Delta,$ since both $dd^{c}\phi_{0}$ and $dd^{c}\phi^{\tau}$ are smooth and strictly positive on $X-\Delta.$ But this means that
$\partial\bar{\partial}\phi^{\tau}(\cdot,\overline{W^{1,0}})=0\,\,\,\text{on $X-\Delta$}$
and since $i\partial\bar{\partial}\phi^{\tau}$ is smooth and strictly positive on $X-\Delta$ it follows that $W=0$ on $X-\Delta.$ Thus, by continuity, $W=0$ on $X,$ which is a contradiction. Hence, 3.23 holds. Finally, the inequality 3.24 is shown in essentially the same way as in the proof of [21, Lemma 9.1]. More precisely, in [21, Lemma 9.1] it is assumed that the corresponding metric $\phi^{\tau}$ is independent of the real part of $\tau$ (in our notation) and that $\phi^{\tau}$ satisfies the complex geodesic equation, that is the local function $\Phi(z,\tau)\coloneqq\phi^{\tau}(z)$ satisfies
(3.25) $(dd^{c}\Phi)^{n+1}=0\,\,\,\text{on $X\times\mathbb{C}$}$
But the proof only uses the latter equation and this equation is satisfied for our $\phi^{\tau}$ (as is well-known and can be seen by locally making the change of variable, $w\coloneqq g_{\tau}(z)).$
@@
Next, let us show that $\mu$ is a critical point of $E$ on the orbit $\mathcal{G}\nu$ iff $\mu\in\mathcal{P}(X)_{0}.$ Since $E$ is $\mathcal{K}$–invariant it will be enough to show that for any real vector field $W$ on $X$ induced by an element in $\mathfrak{k}\subset\mathfrak{g},$
(3.26) $\frac{d}{dt}E(\mu_{t})_{|t=0}=-\int h_{W}\mu,\,\,\,\>\mu_{t}\coloneqq\exp(tJW)^{*}\mu$
@@
where $h_{W}$ is the Hamiltonian function corresponding to $W$ and $\omega_{0}.$ First recall that $h_{W}\coloneqq\mathcal{L}_{JW}\phi_{0},$ (see Section 3.2.8). To prove formula 3.26 note that, by 3.10, denoting by $\phi$ the potential of $\mu,$
$\frac{d}{dt}E(\mu_{t})|_{t=0}=-\int(\phi-\phi_{0})\frac{d\mu_{t}}{dt}|{}_{t=0}=-\int(\phi-\phi_{0})\mathcal{L}\mu,$
Integrating by parts gives
$\frac{d}{dt}E(\mu_{t})|_{t=0}=\int\mathcal{L}(\phi-\phi_{0})\mu=\int\mathcal{L}\phi\frac{(dd^{c}\phi)^{n}}{V}-\int h_{W}\mu,$
which proves 3.26, using the previous lemma and the assumption that the Futaki character of $(X,\Delta)$ vanishes.
Finally, take $\mu\in\mathcal{P}(X)_{0}.$ If $E(\mu)=\infty,$ then $E(g^{*}\mu)=\infty$ for all $g\in\mathcal{G}.$ Hence, we may assume that $E(\mu)<\infty,$ which means that $E(g^{*}\mu)<\infty$ for all $g\in\mathcal{G}.$ Consider, as before, the function $g\mapsto E(g^{*}\mu)$ on $\mathcal{G}/\mathcal{K}.$ It follows from the inequality $\partial_{\tau}\bar{\partial}_{\tau}E(g_{{}_{\tau}}^{*}\nu)\geq 0,$ shown above when $\nu$ is a volume form, and a standard approximation argument, that $E(g^{*}\mu)$ is convex on the symmetric space $\mathcal{G}/\mathcal{K}.$ Moreover, as before, $\frac{d}{dt}E(\mu_{t})_{|t=0}=0$ using the previous lemma, that applies since $E(\mu)<\infty$ (which equivalently means that $\mathcal{E}(\phi_{\mu})>-\infty).$ Since any critical point of a convex function $f$ on the symmetric space $\mathcal{G}/\mathcal{K}$ minimizes $f$ this shows that $\mu$ minimizes $E$ on $\mathcal{G}\mu,$ as desired. ∎
@@
By the previous proposition, the following map is well-defined:
$\pi_{0}\colon\,\mathcal{V}(X)\rightarrow\mathcal{V}(X)_{0}/\mathcal{K},\,\,\,E\left(\pi_{0}(\nu)\right)=\inf_{\mathcal{G}\nu}E$
where $\mathcal{V}(X)$ denotes the subspace of $\mathcal{P}(X)$ consisting of volume forms, i.e. measures with smooth and strictly positive density and $\mathcal{V}(X)_{0}\coloneqq\mathcal{P}(X)_{0}\cap\mathcal{V}(X).$ More generally, $\pi_{0}$ is well-defined for any volume form $\mu$ in $\mathcal{P}(X)$ with divisorial singularities.
@@
(3.27) $\inf_{\mathcal{V}(X)_{0}}F_{-\gamma}=\inf_{\mathcal{\mathcal{V}}(X)}F_{-\gamma}=\inf_{\mathcal{P}(X)_{0}}F_{-\gamma}=\inf_{\mathcal{P}(X)}F_{-\gamma}$
@@
First note that $F_{-\gamma}$ is $\mathcal{K}$–invariant, since $\phi_{0}$ is $\mathcal{K}$–invariant. Moreover, by definition, $F_{-\gamma}=(1-\gamma)E+F_{-1}.$ Since $F_{-1}$ is $\mathcal{G}$–invariant, under the assumption about the vanishing of the Futaki character, the first statement of the lemma thus follows and, as a consequence, also the first identity in formula 3.27. Thus, to prove the rest of the identities 3.27, it will be enough to show that for any given $\mu$ in $\mathcal{P}(X)$ such that $F_{\gamma}(\mu)<\infty$, there exists $\nu_{j}\in\mathcal{V}(X)$ such that $F(\nu_{j})\rightarrow F(\mu).$ But this follows from [27, Lemma 3.1], implying that for any given $\mu$ such that $E(\mu)<\infty$ and $D(\mu|\mu_{0})<\infty,$ there exists a sequence $\mu_{j}$ converging towards $\mu$ in $\mathcal{P}(X)$ such that $E(\mu_{j})$ and $D(\mu_{j}|\mu_{0})$ converge towards $E(\mu)$ and $D(\mu|\mu_{0}),$ respectively. ∎
@@
showing that $\left|E_{\phi_{0}}(\mu)-E_{\phi_{1}}(\mu)\right|\leq C+\left\|u\right\|_{L^{\infty}},$ as desired. Finally, the second statement follows from combining formulae 3.13 and 3.11 with proposition 3.4 and using that $F_{-1}$ is $\mathcal{G}$–invariant (Section 3.2.7). ∎
We are now ready for the proof of the following result, where the equivalence $1\iff 2$ follows directly from [24, 48]. The equivalence $1\iff 4$ provides partial evidence for the Large Deviation Principle (LDP) in Conjecture A, stated in Section 6.2, as it establishes the lower semicontinuity of the rate functional – a property built into the definition of an LDP.
@@
where $d_{1}$ is the $L^{1}$–Finsler metric on $\mathcal{H}(X,\omega_{0})$, which, on $\{\mathcal{E}(u)=0\}$ satisfies
for some positive constants $A$ and $B.$ Indeed, by 3.11, $E(\frac{\omega_{u}^{n}}{V})=(\mathcal{I}-\mathcal{J})(u)$, which is comparable to $\mathcal{J}(u)$ [10, Lemma 2.13] and by [48, Prop 5.5] the inequalities 3.30 do hold when $E(\frac{\omega_{u}^{n}}{V})$ is replaced by $\mathcal{J}(u).$ This proves $1\iff 2$.
Next, note that, $3\implies 2$ follows from the previous proposition. To prove that $2\implies 3$ we recall that it was pointed out in [48, Remark 3.9] that, by density, the inequality 3.29 is equivalent to the corresponding inequality on the metric completion of the metric space $\left(\mathcal{H}(X,\omega_{0}),d_{1}\right),$ which, by [46], may be identified with $\mathcal{E}^{1}(X,\omega_{0}).$ Combined with the inequalities 3.30 this shows that (after perhaps decreasing $\epsilon$ and increasing $C$) there exist $\epsilon,C>0$ such that
@@
In particular, restricting this inequality to $\mathcal{P}(X)_{0},$ and using the last statement of Prop 3.4, yields $F(\mu)\geq\epsilon E(\mu)-C$ on $\mathcal{P}(X)\cap\{E<\infty\},$ i.e. $F_{-(1+\epsilon)}\geq-C$ on $\mathcal{P}(X)\cap\{E<\infty\}$ and hence, by definition, on $\mathcal{P}(X),$ as desired. To prove that $1\iff 4,$ first assume that $F_{-1}$ is lsc on $\mathcal{P}(X)_{0}.$ Since $\mathcal{P}(X)_{0}$ is compact it follows that $F_{-1}$ admits a minimizer in $\mathcal{P}(X)_{0}.$ Such a minimizer is KE, since the Futaki character vanishes (see below). Conversely, assume that $(X,\Delta)$ admits a KE metric. Take a sequence $\mu_{j}$ converging towards $\mu_{\infty}$ in $\mathcal{P}(X)_{0}.$ To prove that $F_{-1}$ is lsc we may as well assume that $F_{-1}(\mu_{j})\leq C.$ Using $1\iff 3,$ it follows that $D(\mu_{j}|\mu_{\phi_{0}})\leq C.$ Thus, by [24, Thm 2.17], $F_{-1}(\mu_{\infty})\leq\liminf F_{-1}(\mu_{j}),$ as desired.
To prove the last statement of the theorem, recall that any KE volume form $\mu$ in $\mathcal{P}(X)$ minimizes $F_{-1}$ on $\mathcal{P}(X)$ [24]. Thus, by Lemma 3.5, there exists a KE volume form $\pi_{0}(\mu)$ minimizing $F_{-1}$ on $\mathcal{P}(X)_{0}$ (but, a priori, $\pi_{0}(\mu)$ is only determined modulo $\mathcal{K}).$ To prove that there exists a unique KE volume form $\mu_{\text{KE}}$ in $\mathcal{P}(X)_{0},$ we recall that, as shown in [21, Lemma 6.3], for any given $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)})$ there exists a unique KE volume form $\mu_{-1}$ in $\mathcal{P}(X)$ minimizing $E$ on the space of KE volume forms in $\mathcal{P}(X),$ i.e. on $\mathcal{G}\mu_{-1}.$ Moreover, $\mu_{-1}$ arises as the limit 3.28. In the present case, where $\phi_{0}$ is assumed $\mathcal{K}$–invariant, it thus follows from Prop 3.4 that $\mu_{-1}\in\mathcal{P}(X)_{0}$ and conversely, any KE volume form in $\mathcal{P}(X)_{0}$ coincides with $\mu_{-1}.$ It also follows from the aforementioned uniqueness that $\mu_{\text{KE}}$ is $\mathcal{K}$–invariant. ∎

### v3

@@ 3.2.9. The moment maps $\boldsymbol{m}$ and $\boldsymbol{m}_{N}$
where $h_{W}$ denotes the Hamiltonian of a real vector field $W$ on $X$ corresponding to an element $W$ in $\mathfrak{k},$ defined by $h_{W}\coloneqq\mathcal{L}_{JW}\phi$ (defined as in formula 3.20). We will denote by $\boldsymbol{m}_{N}$ the moment map of the diagonal action of the group $\mathcal{K}$ on $X^{N},$ endowed with the symplectic form
@@
Moreover, 0 is a regular value of $\boldsymbol{m}_{N}$ iff all the stabilizers in $\mathcal{K}$ of the points in $\left\{\boldsymbol{m}_{N}=0\right\}$ are finite iff the semistable locus of $X^{N}$ coincides with the polystable locus [83, Cor 3.10].
@@ 3.3. The structure of the space $\mathcal{P}(X)_{0}$ of probability measures with vanishing moment
Assume that the Futaki character of $(X,\Delta)$ vanishes. Then, given $\nu\in\mathcal{P}(X)$ such that $E(\nu)<\infty,$ there exists $\mu\in\mathcal{G}\nu$ minimizing $E$ on $\mathcal{G}\nu.$ Moreover, $\mu$ minimizes $E$ on $\mathcal{G}\nu$ iff $\mu\in\mathcal{G}\nu\cap\mathcal{P}(X)_{0}$ and such a measure $\mu$ is uniquely determined modulo the action of $\mathcal{K}.$
@@
First note that $E$ is $\mathcal{K}$–invariant (since $\omega_{0}$ is $\mathcal{K}$–invariant). Given $\nu\in\mathcal{P}(X)$ we thus obtain a function on the quotient $\mathcal{K}\backslash\mathcal{G}$ of $\mathcal{G}$ by the left action of $\mathcal{K}:$
(3.25) $\mathcal{K}\backslash\mathcal{G}\rightarrow\mathbb{R},\,\,\,f(g):=E(g_{*}\nu),$
@@
Since $\mathcal{G}=\mathcal{K}\exp(J\mathfrak{k})$ (by polar decomposition of complex reductive Lie groups) the function $f$ can be identified with a function on $\exp(J\mathfrak{k}).$ Next, let us show that, given $\nu\in\mathcal{P}(X)$ such that $E(\nu)<\infty,$ $\mu$ is a critical point of $E$ on the orbit $\mathcal{G}\nu$ iff $\mu\in\mathcal{P}(X)_{0}.$ Since $E$ is $\mathcal{K}$–invariant it will be enough to show that for any real vector field $W$ on $X$ induced by an element in $\mathfrak{k}\subset\mathfrak{g}$ the following formula holds for any given $\mu\in\mathcal{P}(X)$ such that $E(\mu)<\infty$:
(3.26) $\left.\frac{df(\exp(tJW))}{dt}\right|_{t=0}\coloneqq\frac{d}{dt}E(\exp(tJW)_{*}\mu)_{|t=0}=\int h_{W}\mu,$
@@
where $h_{W}$ is the Hamiltonian function corresponding to $W$ and $\omega_{0}.$ We start by proving the formula when $\mu$ is a smooth form. Set $\mu_{t}\coloneqq\exp(tJW)_{*}\mu.$ By 3.10, denoting by $\phi$ the potential of $\mu,$
$\frac{d}{dt}E(\mu_{t})|_{t=0}=-\int(\phi-\phi_{0})\frac{d\mu_{t}}{dt}|{}_{t=0}=\int(\phi-\phi_{0})\mathcal{L}_{JW}\mu,$
Integrating by parts and using that $h_{W}\coloneqq\mathcal{L}_{JW}\phi_{0}$ (see Section 3.2.9) gives
$\frac{d}{dt}E(\mu_{t})|_{t=0}=-\int\mathcal{L}_{JW}(\phi-\phi_{0})\mu=-\int\mathcal{L}_{JW}\phi\frac{(dd^{c}\phi)^{n}}{V}+\int h_{W}\mu,$
which proves 3.26, using the previous lemma and the assumption that the Futaki character of $(X,\Delta)$ vanishes. Now assume only that $E(\mu)<\infty$ and take a sequence $u_{j}$ in $\mathcal{H}(X,\omega_{0})$ decreasing to the potential $u_{\mu}$ of $\mu.$ Since $u_{\mu}\in\mathcal{E}^{1}(X,\omega)$ it follows that $\mu_{j}:=\omega_{u_{j}}^{n}/V\rightarrow\mu$ and $E(\mu_{j})\rightarrow E(\mu),$ by [25, Prop 1.1]. Likewise, given an automorphism $g$ of $X,$ the sequence $g\cdot u_{j}$ in $\mathcal{H}(X,\omega_{0})$ decreases to $g\cdot u_{\mu}$ and thus $g_{*}\mu_{j}\rightarrow g_{*}\mu$ and $E(g_{*}\mu_{j})\rightarrow E(g_{*}\mu)$. Now, applying the previous case to $\mu_{j}$ and $tW$ for all $t\in]0,\epsilon]$ gives
$\epsilon^{-1}\left(E(\exp(\epsilon JW)_{*}\mu_{j})-E(\mu_{j})\right)=\epsilon^{-1}\int_{0}^{\epsilon}\left(\int_{X}h_{tW}dt\right)\mu_{j}dt.$
Letting $j\rightarrow\infty$ and using that that $h_{tW}$ is continuous thus gives
$\epsilon^{-1}\left(E(\exp(\epsilon JW)_{*}\mu)-E(\mu)\right)=\epsilon^{-1}\int_{X}\left(\int_{0}^{\epsilon}h_{tW}dt\right)\mu dt.$
Finally, letting $\epsilon\rightarrow 0$ and using that $h_{tW}(x)$ is continuous wrt $(t,x)$ proves formula 3.26.
Now assume that $\nu\in\mathcal{P}(X)$ and $E(\nu)<\infty.$ Then $E(g_{*}\nu)<\infty$ for all $g\in\mathcal{G}$ and we will next show that the function $f,$ defined by formula, is strictly convex, when identified with a function on the real vector space $\exp(J\mathfrak{k}).$ To this end fix a real vector field $W$ on $X$ induced by an element in $\mathfrak{k}\subset\mathfrak{g}.$ Differentiating formula 3.26 gives
$\frac{d^{2}}{d^{2}t}_{|t=0}E(\exp(tJW)_{*}\nu)=\int_{X}dH(JW)\nu=\int_{X}\omega_{0}(W,JW)\nu,$
using that $h$ is a Hamiltonian function for $W.$ Since $\omega_{0}(W,JW)$ is non-negative, applying the previous formula to any element in $\exp(\mathbb{R}JW)_{*}\nu$ shows that $t\mapsto E(\exp(tJW)_{*}\nu)$ is a smooth convex function, which is, in fact, strictly convex. Indeed, $\omega(W,JW)$ only vanishes on the subvariety of $X$ defined by the zero-locus of $V$ and $\nu$ does not put mass on analytic subvarieties (since $E(\nu)<\infty,$ $\nu$ does not charge pluripolar subsets [25]).
It follows directly from the strictly convexity of $f$ that its critical points are unique. All that remains is thus to prove that $f$ is proper on $\mathcal{K}/\mathcal{G}$ (and thus admits a critical point). First assume that $\nu=\omega_{0}^{n}/V,$ which minimizes $E$ on all of $\mathcal{P}(X).$ Then $f$ is a strictly convex function admitting a minimum (when $g$ is the identity), which implies that $f$ is proper. Next, for any given $\nu\in\mathcal{P}(X)$ such that $E(\nu)<\infty$ there exists a unique $u\in\mathcal{E}^{1}(X,\omega)$ such that $\omega_{u}^{n}/V=\nu$ and $\mathcal{E}(u)=0.$ Since we have assumed that the Futaki character of $(X,\Delta)$ vanishes, it follows that $\mathcal{E}(u\cdot g)=0$ for all $g\in\mathcal{G},$ using Lemma 3.3. Hence, using the inequality 3.14, it is equivalent to show that the function $g\mapsto d_{1}(u\cdot g,0)$ is proper. Having already shown that $g\mapsto E(g^{*}\omega_{0}^{n}/V)$ is proper – which is equivalent to $d_{1}(0\cdot g,0)$ being proper – it will thus be enough to show that $d_{1}(u\cdot g,0)-d_{1}(0\cdot g,0)$ is bounded by a constant independent of $g.$ But, applying the triangle inequality to $d_{1}$ yields
$\left|d_{1}(u\cdot g,0)-d_{1}(0\cdot g,0)\right|\leq d_{1}(u\cdot g,0\cdot g)=d(u,0),$
using, in the equality, that $\mathcal{G}$ preserves the metric $d_{1}$ [61, Lemma 5.9]. ∎
@@
By the previous proposition, the following map is well-defined, if the Futaki character of $(X,\Delta)$ vanishes:
$\pi_{0}\colon\,\mathcal{P}(X)\cap\{E<\infty\}\rightarrow\mathcal{P}(X)_{0}\cap\{E<\infty\}/\mathcal{K},\,\,\,E\left(\pi_{0}(\nu)\right)=\inf_{\mathcal{G}\nu}E$
@@
(3.27) $\inf_{\mathcal{P}(X)_{0}}F_{-\gamma}=\inf_{\mathcal{P}(X)}F_{-\gamma}$
@@
First note that $F_{-\gamma}$ is $\mathcal{K}$–invariant, since $\phi_{0}$ is $\mathcal{K}$–invariant. Moreover, by definition, $F_{-\gamma}=(1-\gamma)E+F_{-1}.$ Since $F_{-1}$ is $\mathcal{G}$–invariant, under the assumption about the vanishing of the Futaki character, the first statement of the lemma thus follows and, as a consequence, also the first identity in formula 3.27. Thus, to prove the rest of the identities 3.27, it will be enough to show that for any given $\mu$ in $\mathcal{P}(X)$ such that $F_{-\gamma}(\mu)<\infty$, there exists $\nu_{j}\in\mathcal{V}(X)$ such that $F(\nu_{j})\rightarrow F(\mu).$ But this follows from [29, Lemma 3.1], implying that for any given $\mu$ such that $E(\mu)<\infty$ and $D(\mu|\mu_{0})<\infty,$ there exists a sequence $\mu_{j}$ converging towards $\mu$ in $\mathcal{P}(X)$ such that $E(\mu_{j})$ and $D(\mu_{j}|\mu_{0})$ converge towards $E(\mu)$ and $D(\mu|\mu_{0}),$ respectively. ∎
@@
showing that $\left|E_{\phi_{0}}(\mu)-E_{\phi_{1}}(\mu)\right|\leq C+\left\|u\right\|_{L^{\infty}},$ as desired. Finally, the second statement follows from combining formulae 3.15 and 3.11 with proposition 3.4 and using that $F_{-1}$ is $\mathcal{G}$–invariant (Section 3.2.8). ∎
We are now ready for the proof of the following result, where the equivalence $1\iff 2$ follows directly from [26, 61]. The equivalence $1\iff 4$ provides partial evidence for the Large Deviation Principle (LDP) in Conjecture A, stated in Section 6.3, as it establishes the lower semicontinuity of the rate functional – a property built into the definition of a LDP.
@@
where $d_{1}$ is the $L^{1}$–Finsler metric on $\mathcal{H}(X,\omega_{0})$ (see Section 3.2.6). Using the inequality 3.14, this proves $1\iff 2$. Next, note that, $3\implies 2$ follows from the previous proposition. To prove that $1\implies 3$ we recall that it was pointed out in [61, Remark 3.9] that, by density, the inequality 3.29 is equivalent to the corresponding inequality on the metric completion of the metric space $\left(\mathcal{H}(X,\omega_{0}),d_{1}\right),$ which, by [59], may be identified with $\mathcal{E}^{1}(X,\omega_{0}).$ Combined with the inequalities 3.14, this shows that, assuming $1,$ there exist $\epsilon,C>0$ such that
@@
In particular, restricting this inequality to $\mathcal{P}(X)_{0},$ and using Prop 3.4, yields $F(\mu)\geq\epsilon E(\mu)-C$ on $\mathcal{P}(X)\cap\{E<\infty\},$ i.e. $F_{-(1+\epsilon)}\geq-C$ on $\mathcal{P}(X)\cap\{E<\infty\}$ and hence, by definition, on $\mathcal{P}(X),$ as desired. To prove the remaining equivalence $1\iff 4,$ first assume that $F_{-1}$ is lsc on $\mathcal{P}(X)_{0}.$ Since $\mathcal{P}(X)_{0}$ is compact it follows that $F_{-1}$ admits a minimizer in $\mathcal{P}(X)_{0}.$ Such a minimizer is KE, since the Futaki character vanishes (see below). Conversely, assume that $(X,\Delta)$ admits a KE metric. Take a sequence $\mu_{j}$ converging towards $\mu_{\infty}$ in $\mathcal{P}(X)_{0}.$ To prove that $F_{-1}$ is lsc we may as well assume that $F_{-1}(\mu_{j})\leq C.$ Using $1\iff 3,$ it follows that $D(\mu_{j}|\mu_{\phi_{0}})\leq C.$ Thus, by [26, Thm 2.17], $F_{-1}(\mu_{\infty})\leq\liminf F_{-1}(\mu_{j}),$ as desired.
To prove the last statement of the theorem, recall that any KE volume form $\mu$ in $\mathcal{P}(X)$ minimizes $F_{-1}$ on $\mathcal{P}(X)$ [26]. Thus, by Lemma 3.5, there exists a KE volume form $\pi_{0}(\mu)$ minimizing $F_{-1}$ on $\mathcal{P}(X)_{0}$ (but, a priori, $\pi_{0}(\mu)$ is only determined modulo $\mathcal{K}).$ To prove that there exists a unique KE volume form $\mu_{\text{KE}}$ in $\mathcal{P}(X)_{0},$ we recall that, as shown in [22, Lemma 6.3], for any given $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)})$ there exists a unique KE volume form $\mu_{-1}$ in $\mathcal{P}(X)$ minimizing $E$ on the space of KE volume forms in $\mathcal{P}(X),$ i.e. on $\mathcal{G}\mu_{-1}.$ Moreover, $\mu_{-1}$ arises as the limit 3.28. In the present case, where $\phi_{0}$ is assumed $\mathcal{K}$–invariant, it thus follows from Prop 3.4 that $\mu_{-1}\in\mathcal{P}(X)_{0}$ and conversely, any KE volume form in $\mathcal{P}(X)_{0}$ coincides with $\mu_{-1}.$ It also follows from the aforementioned uniqueness that $\mu_{\text{KE}}$ is $\mathcal{K}$–invariant. ∎